OpenStudy (trumpdonald):

HELPPP MEMEMEME!!!! What is the equation of the line x–3y=18 in slope-intercept form? Drag and drop the appropriate number, symbol, or variable to each box. X, Y, -18, -9, -3, -1, 1/3,6 , 9 ,18, +, -, =.

OpenStudy (trumpdonald):

HELP!!!!

Vocaloid (vocaloid):

start by subtracting x from both sides what do you get?

Vocaloid (vocaloid):

still there?

OpenStudy (trumpdonald):

idk

OpenStudy (idkwut):

Try it.

OpenStudy (trumpdonald):

i dont know how to

OpenStudy (idkwut):

You have to turn the equation into y= form. It's basic algebra from there

OpenStudy (trumpdonald):

subtract x

OpenStudy (trumpdonald):

how?

OpenStudy (idkwut):

OK good. x-3y = 18 x-x - 3y = 18 - x 3y = 18 - x Take it from here.

OpenStudy (idkwut):

How do we get y by itself?

OpenStudy (trumpdonald):

um.....we take away x?

OpenStudy (idkwut):

Slope intercept formula is y = mx + b

OpenStudy (idkwut):

You now need to get Y by itself.

OpenStudy (trumpdonald):

do you subtract y like x?

OpenStudy (idkwut):

When a number is next to a variable, it is being multiplied. 3y is the same as saying 3 * Y

OpenStudy (trumpdonald):

ya but what is y i dont get how to subtract or mulitply stuff like that

OpenStudy (idkwut):

What is the opposite of multiplication?

OpenStudy (trumpdonald):

dividing

OpenStudy (idkwut):

yes. so now you basically just divide both sides by 3.

OpenStudy (idkwut):

slope-intercept form: y = mx + b where m = the slope and b = the y-intercept

OpenStudy (trumpdonald):

i still dont know whatd going on

OpenStudy (idkwut):

starting from where?

OpenStudy (trumpdonald):

give meh the answeeeerrrr

OpenStudy (trumpdonald):

:)

OpenStudy (idkwut):

No, you were doing good.

OpenStudy (trumpdonald):

no i dont get it

OpenStudy (idkwut):